OpenStudy (anonymous):

Evaluate the logarithm simplify log2 2^x

OpenStudy (anonymous):

\[\log_{2} 2^{x}\]

OpenStudy (kittiwitti1):

as @jdoe0001 said...\[\log_{b}a=x\rightarrow a^{x}=b\]just input the variable values... and solve :-)

OpenStudy (anonymous):

i dont understand that

OpenStudy (jdoe0001):

well, same cancellation rule :)

OpenStudy (jdoe0001):

hmmm actually, hold the mayo, is a cancellation rule, just not that one =)

OpenStudy (anonymous):

so is the answer just 2x

OpenStudy (anonymous):

or x

OpenStudy (jdoe0001):

\(\bf log_{\color{red}{ a}}{\color{red}{ a}}^x\implies x\qquad \qquad thus \\ \quad \\ log_{\color{red}{ 2}}({\color{red}{ 2}}^x)\implies ?\)

OpenStudy (anonymous):

so the answer is x?

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

thanks :)

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

i have another

OpenStudy (anonymous):

can u help

OpenStudy (anonymous):

use \[\log_{8} 9 \approx 1.057 and \log_{8} 4 \approx 0.667 \]

OpenStudy (anonymous):

to evaluate

OpenStudy (anonymous):

\[\log_{8 } \frac{ 9 }{ 4 }\]

OpenStudy (kittiwitti1):

the equation should be written in this format...\[\frac{\log_{3}9}{\log_{3}4}\]now you have those two values... just put them into this equation and simplify...

OpenStudy (anonymous):

where did the 3 come from?

OpenStudy (kittiwitti1):

\[\frac{\log_{8}9}{\log_{8}4}\rightarrow\frac{1.057}{0.667}\]sorry, base is 8, screen is very small

OpenStudy (kittiwitti1):

i misread... apologies haha

OpenStudy (anonymous):

its all good :)

OpenStudy (anonymous):

i got 1.58

OpenStudy (kittiwitti1):

does it say how many significant figures to round up/down to...?

OpenStudy (kittiwitti1):

or is that the full answer... :-o

OpenStudy (anonymous):

well 1.585 was the answer choice

OpenStudy (anonymous):

i have another thats more difficult like that one i think

OpenStudy (kittiwitti1):

yes well... make sure you have the same amount of digits in your answer when rounding, as the answer options... okay

OpenStudy (kittiwitti1):

i think you should make a new post though... this post is very long and laggy lol

OpenStudy (anonymous):

use \[\log_{3} 8\approx 1.893 and \log_{3} 5\approx 1.465 \]

OpenStudy (anonymous):

to evaluate

OpenStudy (anonymous):

\[\log_{3} 40\]

OpenStudy (kittiwitti1):

oh... this is one of those equations... so basically they are saying they got it like this:\[\log_{3}8+\log_{3}5\rightarrow\log_{3}(8\times5)\rightarrow\log_{3}40\]you get it yes?

OpenStudy (kittiwitti1):

so they give you values for the 8 and 5... you just need to put in those equations and add them together... \[\log_{3}8+\log{3}5\rightarrow 1.893+1.465=?\]

OpenStudy (kittiwitti1):

do you get it?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

can you help with another

OpenStudy (kittiwitti1):

ok good... now here is a set of the log rules so you understand why i am doing it this way... http://www.rapidtables.com/math/algebra/Logarithm.htm

OpenStudy (kittiwitti1):

it will help you solve for the rest of the equations with this kind of stuff... i cannot help you on every single problem, you will not learn :-P

OpenStudy (anonymous):

yeah but like they are all different lol

OpenStudy (kittiwitti1):

well... if you get the basic rules in the link i sent you... you should be good :-)

OpenStudy (kittiwitti1):

here is a really good video link if you are confused on how to apply the formulas... https://www.khanacademy.org/math/algebra2/exponential-and-logarithmic-functions/properties-of-logarithms/v/using-multiple-logarithm-properties-to-simplify :-)

OpenStudy (anonymous):

okay thanks

OpenStudy (kittiwitti1):

np... good luck ♣